OpenStudy (anonymous):

A grocer is dividing a 24 lb box of green beans in to packages. How many 1 1/2 lb packages of green beans can he make? A recipe calls for 1 3/4 c of flour. If you make 3 1/2 times the recipe how much flour will you need

OpenStudy (whpalmer4):

Dividing 24 pounds of green beans into 1 1/2 pound packages: \[\frac{24}{1\frac{1}{2}} =\] Making a bigger batch: \[1\frac{3}{4}*3\frac{1}{2} = \] For both problems, I would start by converting the mixed fractions to improper fractions (numerator larger than denominator).